# LEP3: A High-Luminosity $e^{+}e^{-}$ Higgs & Electroweak Factory in the LHC Tunnel

### v1

@@ 4.1 Higgs Boson Physics
Table 2: Comparison of precision measurements of Higgs boson parameters. The HL-LHC numbers are taken from the ATLAS+CMS submission to the ESPP26, the LEP3 numbers are scaled from the FCC-ee numbers in [7] (the statistical part of the error is scaled up by $\sim 2.6$).
@@ 4.2 Electroweak and Flavour Physics
Table 3: Experimental precision (statistical + systematic) of a selection of electroweak measurements at LEP3 compared with the present world-average precisions.
@@
With $2\times 10^{12}$ Z boson decays recorded, we expect $2.5\times 10^{11}B^{0}/\bar{B}^{0},2.5\times 10^{11}B^{+}/B^{-},6\times 10^{10}B_{s}^{0},1.6\times 10^{9}B_{c}$ decays, as well as $2\times 10^{12}$ charm quark pairs and $7\times 10^{10}\tau^{+}\tau^{-}$ pairs. This enables precision and extensive flavour physics programme at LEP3, largely complementary to the HL-LHC and Belle II programs. In particular, precision measurements of fundamental parameters of tau leptons (lifetime, mass, leptonic branching fraction), rare $B_{s}^{0}$ and $B_{c}$ decays and b baryon decays with missing energy, and charm physics will yield very rich and often unique results.
@@
As has been discussed, e.g., in [8], precision electroweak measurements around the Z peak provide indirect sensitivity to new physics at scales tens of TeV scale through higher-order loop effects. These indirect effects provide probes of potential deviations from the Standard Model that are complementary to those obtainable from measurements at higher energies above the Z pole. Explicit examples of how accuracy complements energy for operators within the SMEFT framework are given in [8]. A high-precision Tera-Z programme may thus anticipate aspects of physics runs at higher energies and provide a wide scope of quantum exploration of the TeV scale. See Table 3for the expected perfomance of LEP3.

### v2

@@ 4.1 Higgs Boson Physics
Table 4: The precision for Higgs boson couplings measurements. The errors for LEP3 have been increased with respect to the recent FCC-ee submission  [9], by a factor $\sqrt{22/4.2}$. In red are HL-LHC values that are unlikely to be improved by $e^{+}e^{-}$ Higgs boson factories.
@@ 4.2 Electroweak and Flavour Physics
Table 5: Experimental precision (statistical + systematic) of a selection of electroweak mea- surements at LEP3. The statistical errors are scaled by a factor $\sqrt{6/1.7}$ from the FCC-ee submission  [10],  [11],  [9] whilst the systematics ones are kept the same.
@@
With $1.7\times 10^{12}$ Z boson decays recorded, we expect $2.1\times 10^{11}B^{0}/\bar{B}^{0},2.1\times 10^{11}B^{+}/B^{-},5.1\times 10^{10}B_{s}^{0},1.3\times 10^{9}B_{c}$ decays, as well as $1.7\times 10^{12}$ charm quark pairs and $6\times 10^{10}\tau^{+}\tau^{-}$ pairs. This enables precision and extensive flavour physics programme at LEP3, largely complementary to the HL-LHC and Belle II programs. In particular, precision measurements of fundamental parameters of tau leptons (lifetime, mass, leptonic branching fraction), rare $B_{s}^{0}$ and $B_{c}$ decays and b baryon decays with missing energy, and charm physics will yield very rich and often unique results.
@@
As has been discussed, e.g., in [12], precision electroweak measurements around the Z peak provide indirect sensitivity to new physics at scales tens of TeV through higher-order loop effects. These indirect effects provide probes of potential deviations from the Standard Model that are complementary to those obtainable from measurements at higher energies above the Z pole. Explicit examples of how accuracy complements energy for operators within the SMEFT framework are given in [12]. A high-precision Tera-Z programme may thus anticipate aspects of physics runs at higher energies and provide a wide scope of quantum exploration of the TeV scale. See Table 5 for the expected perfomance of LEP3.